# APTx Neuron: A Unified Trainable Neuron Architecture Integrating Activation and Computation

### v4

@@ Abstract
We propose the APTx Neuron, a novel, unified neural computation unit that integrates non-linear activation and linear transformation into a single trainable expression. The APTx Neuron is derived from the APTx activation function, thereby eliminating the need for separate activation layers and making the architecture both computationally efficient and elegant. The proposed neuron follows the functional form $y=\sum_{i=1}^{n}((\alpha_{i}+\tanh(\beta_{i}x_{i}))\cdot\gamma_{i}x_{i})+\delta$, where all parameters $\alpha_{i}$, $\beta_{i}$, $\gamma_{i}$, and $\delta$ are trainable. We validate our APTx Neuron-based architecture on the MNIST dataset, achieving up to $96.69\%$ test accuracy within 11 epochs using approximately 332K trainable parameters. The results highlight the superior expressiveness and computational efficiency of the APTx Neuron compared to traditional neurons, pointing toward a new paradigm in unified neuron design and the architectures built upon it. Source code is available at https://github.com/mr-ravin/aptx_neuron.
@@ 1 Introduction
Custom activation functions such as Swish [15], Mish [11], and ELU [2] have demonstrated superior performance compared to the traditional ReLU [12] in various deep learning applications. These advancements reflect a broader trend toward more adaptive and expressive non-linearities.
Among recent innovations, the APTx activation function [5, 6] is notable for its parametric, trainable formulation, which can approximate multiple activation behaviors, including Swish and Mish, and can also resemble Tanh-like curves under certain parameter settings. As noted in recent surveys on trainable activation functions [1], such flexibility enables neural networks to better adapt to task-specific requirements during training. The mathematical formulation of the APTx activation function is given in Equation 1.
@@ 2 Background and Motivation
In standard feedforward neural networks, a neuron performs a two-step process: it computes a weighted sum of inputs followed by the application of a non-linear activation function [16]. Mathematically, this is often written as shown in Equation 2.
@@
The APTx activation function [5, 6], due to its parametric nature, already offers adaptive behavior and a rich expressiveness to simulate or interpolate between several standard non-linearities. It introduces trainable parameters $\alpha$, $\beta$, and $\gamma$ into the activation itself, allowing it to learn optimal non-linear transformations during training. This results in improved flexibility and performance across tasks and architectures.
@@ 3 APTx Neuron
The APTx Neuron is a novel computational unit that unifies linear transformation and non-linear activation into a single, expressive formulation. Inspired by the parametric APTx activation function, this neuron architecture removes the strict separation between computation and activation, allowing both to be learned as a cohesive entity. It is designed to enhance representational flexibility while reducing architectural redundancy. A key theoretical property of the APTx Neuron is that it naturally retains the universal approximation capability [4].
@@ 3.2 Relationship to Traditional Neurons and Activations
    The form presented in Equation 6 is structurally identical to a conventional linear neuron with learnable weights and bias. Consequently, the universal approximation property [4] is inherently preserved in the APTx Neuron.
  • Pure Activation Behavior: When $\delta=0$ and input is passed as a fixed vector (e.g., $x=[x_{1},x_{2},...,x_{n}]$), the APTx Neuron computes a non-linear transformation akin to a composite APTx activation function:
@@ 3.4 Parameter Overhead and Efficiency
Importantly, the formulation of the APTx Neuron preserves the universal approximation capability [4] of neural networks. By embedding trainable, input-wise non-linearities directly within each neuron, the APTx Neuron goes beyond traditional designs where approximation power depends on stacking fixed activations atop linear transformations. Each input dimension in the APTx Neuron can independently learn its own transformation behavior, both linear and non-linear, enabling more efficient, compact, and expressive modeling, even in shallow architectures.
@@ 4 Architecture, Training, and Results
To evaluate the effectiveness of the proposed APTx Neuron, we implemented a custom fully connected feedforward neural network using APTx Neurons in PyTorch [13]. This section outlines the design, training pipeline, and performance metrics on the MNIST dataset [3].
@@ 4.1 Neural Network Design
The source code for the APTx Neuron-based architecture and the MNIST [3] experiment, implemented in Python [14] using the PyTorch library [13], is available at our GitHub repository [9]. In addition, a pip-installable Python package implementing the APTx Neuron is available on both GitHub [7] and PyPI [8]. It can be installed using the following command: pip install aptx_neuron.
@@ 5 Toward Integration in CNNs and Transformers
The APTx Neuron, as introduced in this paper, is a general-purpose, unified computational unit that combines both activation and transformation into a single trainable expression. Although we have demonstrated its effectiveness within a fully connected feedforward neural network, the design is not limited to MLPs. As a fundamental building block, the APTx Neuron can be readily extended to modern architectures such as convolutional neural networks (CNNs) and transformers.
In CNNs [10], activation functions like ReLU or Swish are typically applied after convolutional filters. By replacing these fixed activations with APTx-style computation, it is possible to create APTx Convolutional Units that learn both spatial filtering and adaptive non-linearity in a unified manner, potentially improving spatial feature learning.
In transformer architectures [17], non-linear transformations are used in the feedforward sublayers and position-wise projections. Substituting those fixed activations with APTx Neurons may allow the model to dynamically adapt its internal representations and activation behavior based on the context or task, an especially valuable trait in attention-based models.
Thus, the APTx Neuron is not only a compact and expressive component for standard networks, but also a promising primitive for redefining core operations in CNNs, Transformers, and hybrid neural systems.
@@ 6 Conclusion
This work introduced the APTx Neuron, a unified, fully trainable neural unit that integrates linear transformation and non-linear activation into a single expression, extending the APTx activation function. By learning per-input parameters $\alpha_{i}$, $\beta_{i}$, and $\gamma_{i}$ for each input $x_{i}$, and a shared bias term $\delta$ within a neuron, the APTx Neuron removes the need for separate activation layers and enables fine-grained input transformation. The APTx Neuron generalizes traditional neurons and activations, offering greater representational power. Our experiments show that a fully connected APTx Neuron-based feedforward neural network achieves $96.69\%$ test accuracy on the MNIST dataset within 11 epochs using approximately 332K trainable parameters, demonstrating rapid convergence and high efficiency. This design lays the groundwork for extending APTx Neurons to CNNs and transformers, paving the way for more compact and adaptive deep learning architectures.

### v5

@@ Abstract
We propose the APTx Neuron, a novel, unified neural computation unit that integrates non-linear activation and linear transformation into a single trainable expression. The APTx Neuron is derived from the APTx activation function, thereby eliminating the need for separate activation layers and making the architecture both optimization-efficient and elegant. The proposed neuron follows the functional form $y=\sum_{i=1}^{n}((\alpha_{i}+\tanh(\beta_{i}x_{i}))\cdot\gamma_{i}x_{i})+\delta$, where all parameters $\alpha_{i}$, $\beta_{i}$, $\gamma_{i}$, and $\delta$ are trainable. We validate our APTx Neuron-based architecture on the MNIST dataset, achieving up to $96.69\%$ test accuracy within 11 epochs using approximately 332K trainable parameters. The results highlight the superior expressiveness and training efficiency of the APTx Neuron compared to traditional neurons, pointing toward a new paradigm in unified neuron design and the architectures built upon it. Source code is available at https://github.com/mr-ravin/aptx_neuron.
@@ 1 Introduction
Custom activation functions such as Swish [ramachandran2017searchingactivationfunctions], Mish [misra2019mish], and ELU [clevert2015fast] have demonstrated superior performance compared to the traditional ReLU [10.5555/3104322.3104425] in various deep learning applications. These advancements reflect a broader trend toward more adaptive and expressive non-linearities.
Among recent innovations, the APTx activation function [kumar2022aptx, aptxactivation2022code] is notable for its parametric, trainable formulation, which can approximate multiple activation behaviors, including Swish and Mish, and can also resemble Tanh-like curves under certain parameter settings. As noted in recent surveys on trainable activation functions [APICELLA202114], such flexibility enables neural networks to better adapt to task-specific requirements during training. The mathematical formulation of the APTx activation function is given in Equation 1.
@@ 2 Background and Motivation
In standard feedforward neural networks, a neuron performs a two-step process: it computes a weighted sum of inputs followed by the application of a non-linear activation function [Szandała2021]. Mathematically, this is often written as shown in Equation 2.
@@
The APTx activation function [kumar2022aptx, aptxactivation2022code], due to its parametric nature, already offers adaptive behavior and a rich expressiveness to simulate or interpolate between several standard non-linearities. It introduces trainable parameters $\alpha$, $\beta$, and $\gamma$ into the activation itself, allowing it to learn optimal non-linear transformations during training. This results in improved flexibility and performance across tasks and architectures.
@@ 3 APTx Neuron
The APTx Neuron is a novel computational unit that unifies linear transformation and non-linear activation into a single, expressive formulation. Inspired by the parametric APTx activation function, this neuron architecture removes the strict separation between computation and activation, allowing both to be learned as a cohesive entity. It is designed to enhance representational flexibility while reducing architectural redundancy. A key theoretical property of the APTx Neuron is that it naturally retains the universal approximation capability [HORNIK1989359].
@@ 3.2 Relationship to Traditional Neurons and Activations
    The form presented in Equation 6 is structurally identical to a conventional linear neuron with learnable weights and bias. Consequently, the universal approximation property [HORNIK1989359] is inherently preserved in the APTx Neuron.
  • Pure Activation Behavior: When $\delta=0$ and input is passed as a fixed vector (e.g., $x=[x_{1},x_{2},...,x_{n}]$), the APTx Neuron computes a non-linear transformation akin to a composite APTx activation function as shown in Equation 7.
@@ 3.4 Parameter Overhead and Efficiency
Importantly, the formulation of the APTx Neuron preserves the universal approximation capability [HORNIK1989359] of neural networks. By embedding trainable, input-wise non-linearities directly within each neuron, the APTx Neuron goes beyond traditional designs where approximation power depends on stacking fixed activations atop linear transformations. Each input dimension in the APTx Neuron can independently learn its own transformation behavior, both linear and non-linear, enabling more efficient, compact, and expressive modeling, even in shallow architectures.
@@ 4 Architecture, Training, and Results
To evaluate the effectiveness of the proposed APTx Neuron [kumar2025aptx_arxiv, kumar2025aptx_ssrn], we implemented a custom fully connected feedforward neural network using APTx Neurons in PyTorch [10.5555/3454287.3455008]. This section outlines the design, training pipeline, and performance metrics on the MNIST dataset [6296535].
@@ 4.1 Neural Network Design
The source code for the APTx Neuron-based architecture and the MNIST [6296535] experiment, implemented in Python [python] using the PyTorch library [10.5555/3454287.3455008], is available at our GitHub repository [aptxneuron2025code]. In addition, a pip-installable Python package implementing the APTx Neuron is available on both GitHub [aptx_neuron_package_github] and PyPI [aptx_neuron_package_pip]. It can be installed using the following command: pip install aptx_neuron.
@@ 4.3 Performance Metrics
Although the APTx Neuron performs more operations per forward pass than a standard linear+activation unit, it demonstrates computational efficiency in terms of optimization behavior: the model converges significantly faster, surpassing $96\%$ test accuracy within only 6 epochs and reaching near-zero training loss by epoch 8.
@@ 5 Toward Integration in CNNs and Transformers
The APTx Neuron, introduced in this work and defined in Equation 4, is a unified computational unit that integrates activation and transformation within a single trainable expression. Although its effectiveness has been demonstrated in fully connected feedforward networks, the formulation is not restricted to MLPs. As a general building block, the APTx Neuron and the APTx Activation Function (Equation 1) can be incorporated into convolutional neural networks (CNNs) and Transformer architectures with minimal structural modification.
5.1 Integration in Convolutional Neural Networks
In convolutional neural networks (CNNs) [726791], activation functions such as ReLU, Swish, or Mish are commonly applied after convolutional filtering. The APTx framework supports two forms of integration: (i) replacing only the activation function, or (ii) forming a unified modulation operator on top of the convolutional responses, analogous to the modulation in Equation 4.
5.1.1 Case 1: Activation-Only Replacement:
A convolutional layer produces feature maps $z_{k}$ as defined in Equation 9.
$z_{k}=w_{k}*x,\qquad z_{k}\in\mathbb{R}^{H\times W}$ (9)
where $C_{\text{out}}$ is the number of output channels of the convolutional layer and $H\times W$ denotes the spatial resolution (height and width) of each feature map. The convolution kernels $w_{k}$ remain unchanged, and the APTx Activation Function from Equation 1 is applied element-wise as shown in Equation 10, where parameters $\alpha_{k},\beta_{k},\gamma_{k}$ are shared across spatial locations.
$y_{k}(i,j)=\bigl(\alpha_{k}+\tanh(\beta_{k}z_{k}(i,j))\bigr)\,\gamma_{k}\,z_{k}(i,j)$ (10)
This produces the output $Y=[y_{1},\dots,y_{C_{\text{out}}}]\in\mathbb{R}^{C_{\text{out}}\times H\times W}$.
5.1.2 Case 2: Unified APTx Convolutional Unit:
A more expressive integration applies an APTx-style modulation, similar to the structure in Equation 4, directly to the convolutional responses as expressed in Equation 11.
$y_{k}(i,j)=\bigl(\alpha_{k}+\tanh(\beta_{k}z_{k}(i,j))\bigr)\,\gamma_{k}\,z_{k}(i,j)+\delta_{k}$ (11)
Depending on the learned parameters, the unit may behave like a dynamic activation or a channel-wise neuron-like modulator. The resulting output is $Y=[y_{1},\dots,y_{C_{\text{out}}}]\in\mathbb{R}^{C_{\text{out}}\times H\times W}$.
5.2 Integration in Transformer Architectures
Transformers [10.5555/3295222.3295349] rely on multi-head self-attention and a position-wise feed-forward network (FFN). Both components can benefit from APTx-based units.
5.2.1 Replacing Q, K, and V Projections:
The projections $Q$, $K$, and $V$ as shown in Equation 12 may be replaced by APTxLayer blocks constructed from APTx Neurons (Equation 4) as shown in Equation 13.
$Q=W_{Q}x,\qquad K=W_{K}x,\qquad V=W_{V}x$ (12)
$Q=\mathrm{APTxLayer}(x),\qquad K=\mathrm{APTxLayer}(x),\qquad V=\mathrm{APTxLayer}(x)$ (13)
Replacing the Feed-Forward Layer with an APTx Neuron Layer:
The standard FFN is defined in Equation 14.
$\mathrm{FFN}(x)=W_{2}\,\phi(W_{1}x+b_{1})+b_{2}$ (14)
Replacing $\phi(\cdot)$ with the APTx activation function as $(\alpha+\tanh(\beta u))\,\gamma u$, and letting $u=W_{1}x+b_{1}$, we obtain Equation 15.
$\mathrm{FFN}_{\mathrm{APTxActivation}}(x)=W_{2}\left((\alpha+\tanh(\beta(W_{1}x+b_{1})))\,\gamma(W_{1}x+b_{1})\right)+b_{2}$ (15)
A full FFN block can also be constructed using two stacked APTxLayers, each composed of APTx Neurons, as shown in Equation 16. Since each APTx Neuron (Equation 4) already combines linear transformation, non-linearity, and bias, two such layers are sufficient to mirror the role of the standard two-layer FFN.
$\mathrm{FFN}_{\mathrm{APTxNeuron}}(x)=\mathrm{APTxLayer}_{2}\!\bigl(\mathrm{APTxLayer}_{1}(x)\bigr)$ (16)
These integration pathways illustrate that APTx-based components form flexible and expressive building blocks for CNNs, Transformers, and hybrid neural architectures, extending the applicability of the unified APTx Neuron beyond multilayer perceptrons.
@@ 6 Conclusion
This work introduced the APTx Neuron, a unified, fully trainable neural unit that integrates linear transformation and non-linear activation into a single expression, extending the APTx activation function. By learning per-input parameters $\alpha_{i}$, $\beta_{i}$, and $\gamma_{i}$ for each input $x_{i}$, and a shared bias term $\delta$ within a neuron, the APTx Neuron removes the need for separate activation layers and enables fine-grained input transformation. The APTx Neuron generalizes traditional neurons and activations, offering greater representational power. Our experiments show that a fully connected APTx Neuron-based feedforward neural network achieves $96.69\%$ test accuracy on the MNIST dataset within 11 epochs using approximately 332K trainable parameters, demonstrating rapid convergence and high optimization efficiency. This design lays the groundwork for extending APTx Neurons to CNNs and transformers, paving the way for more compact and adaptive deep learning architectures.